OpenStudy (cheesecakekitten):

@AloneS Solve x^2 - 8x + 4 = 0 A. {-4 ± 26} B. {4 ± 3} C. {-4 ± 212} D. {-4 ± 6}

rebeccaxhawaii (rebeccaxhawaii):

any ideas?

alones (alones):

@CheesecakeKitten @CheesecakeKitten @CheesecakeKitten @CheesecakeKitten YOU THERE? OR AM I ALONE

jhonyy9 (jhonyy9):

do you can calculi the discriminant ?

jhonyy9 (jhonyy9):

x^2 -8x +4 =0 D = b^2 -4ac = ?

jhonyy9 (jhonyy9):

b=-8 a=1 c= 4

jhonyy9 (jhonyy9):

collaborate please !!!